OpenStudy (anonymous):

Use integration by parts to solve the integral of (e^x)(Cos(2x))

OpenStudy (anonymous):

\[\int\limits_{}^{} e ^{x}.\cos 2x\]

OpenStudy (anonymous):

Do you know about Integration by parts??

OpenStudy (anonymous):

\[\int\limits uv = u \int\limits v - \int\limits (u' \times \int\limits v)\]

OpenStudy (anonymous):

yes I know the formula but both functions e^x and cos2x are both decaying functions when you derive or intergrate them. My guess is you'd have to use trig identities on cos2x

OpenStudy (cwrw238):

INT udv = uv - INT vdu where u and v are functions

OpenStudy (anonymous):

See here u = cos(2x) and v = \(e^x\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you use it in the formula???

OpenStudy (anonymous):

Firstly let: \[\int\limits e^x. \cos2x .dx = I\]

OpenStudy (anonymous):

Now apply the formula..

OpenStudy (anonymous):

yes you end up with. \[e ^{x}.-2\sin (2x) - \int\limits \left[ -2\sin (2x) . e ^{x} \right]\]

OpenStudy (anonymous):

where you still have an integral function..

OpenStudy (anonymous):

Where from - come from??

OpenStudy (anonymous):

sorry. I did my integral wrong

OpenStudy (anonymous):

Firstly tell me what is the integral of cos2x??

OpenStudy (anonymous):

(sin2x)/2

OpenStudy (anonymous):

Yes now write it properly in your notebook..

OpenStudy (anonymous):

yes and you're still left with an integral function.

OpenStudy (anonymous):

\[I = e ^{x}.\frac{\sin (2x)}{2} - \int\limits\limits\limits \left[e^x. \frac{\sin (2x)}{2}dx\right]\]

OpenStudy (anonymous):

In the case we have one exponential and one trigonometric function in the Integral we have to do integration twice. One time you have done this.. Now apply the integration to the last term again..

OpenStudy (anonymous):

take the 2 out of the integral in last term..

OpenStudy (anonymous):

do you mean take out 1/2?

OpenStudy (anonymous):

Solve for the last term: \[\int\limits e^x.\sin2xdx\] u = sin2x and v = \(e^x\) Again take the integral.. Doing or not??

OpenStudy (anonymous):

Yes I mean that take 1/2 out of the integral..

OpenStudy (anonymous):

okay um give me a sec

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

is it \[\sin 2x.e ^{x} - \int\limits \cos2x-e ^{x}\]

OpenStudy (anonymous):

where you derive sin2x and integrate e^x

OpenStudy (anonymous):

Can you find your mistakes??

OpenStudy (anonymous):

integral of sin2x is : \(\large \frac{-cos2x}{2}\)

OpenStudy (anonymous):

\[\sin2x.e ^{x} - \int\limits \cos2x.e ^{x}\]

OpenStudy (anonymous):

oh, i derived sin2x not integrated it.

OpenStudy (anonymous):

In first term you have to take integral of sin2x as you did in second..

OpenStudy (anonymous):

i think you meant to say v = sin2x and u = ex instead of the other way around

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

i'm confused, in the second one was i meant to derive or integrate sin2x?

OpenStudy (anonymous):

SOrry..

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We are doing it wrong..

OpenStudy (anonymous):

Let me show the steps just wait..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):